# Intent-Based Access Control: Using LLMs to Intelligently Manage Access Control

### v1

@@ 1. Introduction
Over the last several years, there has been a proliferation of easy-to-use software tools for data science. Organizations of all sizes and from all industries are applying data-driven models to make decisions. This trend, which is often called the democratization of data science (Chou et al., 2014), is as worrying as it is exciting. On one hand, some of humanity’s most difficult scientific challenges are being addressed with data science from particle physics to drug discovery. On the other hand, there are numerous cautionary tales of data misuse, data breaches, data quality issues, and improper data sharing (Monroe-White et al., 2021; Janssen et al., 2020; Kerber, 2020; Seh et al., 2020; Fahey, 2021). Data governance, the process of managing the availability, usability, integrity, and security of the data in enterprise systems, has consequently emerged as a key discipline in enterprise data management. Access control is one of the most important technical challenges in data governance, where a system must ensure that only authorized users can access particular data assets.
In any sizeable organization, it is impractical to exhaustively map individual users to data assets. Thus, role-based access control (RBAC) was a key innovation in data governance and is widely supported by data management platforms where groups of users can be selectively given access (Sandhu et al., [n.d.]). Such forms of access control are well-studied in enterprise SQL databases, where roles, views, and GRANT statements are already used to restrict user access to data (Chaudhuri et al., 2011; Bertino et al., 2011; Bertino and Sandhu, 2005). Similarly, many databases additionally support discretionary access control (DAC) where roles can be given delegation privileges. RBAC and DAC help reduce the administrative burden in defining access control rules for a database.
@@
  • LLM4AC: A system built from a reference architecture for an IBAC-DB implementation that provides a novel format for specifying access control policies, automatically implementing them, and comparing implementations to policies for correctness.
@@ 2.4. Existing NL2SQL Capabilities
Based on these examples and initial experiments, rather than trying to train NL2SQL systems to 100%, we hypothesize if access control policies can be formatted to clearly reflect a role, privilege, and view, and we can correctly represent access control policies in terms of the database roles, views, and privileges that correspond to the policy, then NL2SQL methods adapted for access control can correctly implement access control policies.
@@ 4.2. The Differencer, In-Depth
LLM4AC At Scale. Using LLM4AC to difference two NLACMs with dimensions $m_{1}\times n_{1}$ and $m_{2}\times n_{2}$ requires $O(m_{min}\times n_{min})$ calls to the LLM for privilege subsumption, in the worst case. This may seem prohibitively expensive for large NLACMs, both in terms of runtime and monetary cost of using commercial LLMs, like ChatGPT.
Firstly, we argue that large NLACMs will not be part of a typical workload for LLM4AC. NLACMs are written by users who want to ensure access control rules are implemented correctly on a small, specific set of roles/users and tables/views. They will likely use differencing to test their policy implementations on a test database containing no roles or users, only base tables.
That said, suppose a user now wants to difference with respect to their production database, which already has many users and view definitions. Then, the NLACM generated from the database may be large, but this does not cause a scalability issue in terms of LLM calls because we ask the LLM to choose from a list of values during role-view mapping instead of checking each pair. If the list of values exceeds the LLM’s context window, we find that chunking the list is an accurate strategy. This can also greatly alleviate the scale problem. For example, ChatGPT has a context window length of 8192 tokens, and a list of 10 view definitions, each roughly 50 tokens long, on average, resulting in only 10 calls of 500 tokens each, with 500 more tokens of output. This costs 5500 tokens. This is cheaper compared to 100 pairwise comparisons, each consisting of at least 100 tokens and 1 token of output (a yes/no token) (10001 tokens).
Now, suppose a user wants to create a large NLACM. For example, they want a NLACM that expresses a policy for their whole production database. LLM4AC will not force users to write it themselves. Instead, because NLACMs can be a mix of NL and SQL, users can specify only the key access control rules in NL, and then LLM4AC will perform the role-view mapping step of differencing to determine which roles and views written in the user’s NLACM already have privileges in the database. LLM4AC will then drop such overlapping roles and views and union this NLACM with the NLACM generated from the database.
5. Evaluation
In this section, we propose a benchmark for evaluating IBAC-DB systems (Section 5.1). We use this benchmark to show that our system, LLM4AC, can accurately synthesize and difference access control policies (Sections 5.2 and 5.3), especially compared to other baselines. Finally, we use this benchmark to determine if there are types of access control policies on which LLM4AC is not as accurate (Section 5.4).
For all experiments, the DBMS is PostgreSQL. We use ChatGPT (gpt-3.5-turbo) as the LLM, and ChatGPT’s APIs for all prompting.
@@ 5.1. IBAC-DB Benchmark Design
We design a benchmark to test IBAC-DB systems that implement access control policies in PostgresQL. The goal of the benchmark is to test how accurately an IBAC-DB system synthesizes a SQL implementation of a NLACM policy, and how accurately an IBAC-DB system can compare two NLACMs. We describe the benchmark design for each of IBAC-DB synthesizers and differencers.
@@ 5.1.1. The Differencer Benchmark
Differencer Evaluation Scenarios. It is essential that IBAC-DB system differencers be able to accurately compare a policy to its implementation for compliance, and compare access control policies to ensure they actually match. To this end, we measure differencing with respect to two scenarios: (i) policy auditing: compare a NLACM containing NL to its implementation on the DB, which is a NLACM containing SQL. (ii) policy comparison: compare two NLACMs containing NL.
Differencer Evaluation Metric. We evaluate differencing with respect to: (i) role-view mapping accuracy: how accurately LLM4AC can find the shared roles and views between the two NLACMs, (ii) privilege subsumption accuracy: how accurately LLM4AC can determine whether the privileges contained in the second NLACM subsumed those of the first.
We use F1 scores to measure role-view mapping and privilege subsumption accuracies. Specifically, for role-view mapping, a true positive is a role-view pair shared by both NLACMs that was detected by LLM4AC’s differencing procedure. A true negative is a role-view pair not shared by both NLACMs that was not detected by LLM4AC’s differencing procedure. A false positive is a role-view pair not shared by both NLACMs that LLM4AC’s differencing procedure incorrectly determines is shared. A false negative is a shared role-view pair that LLM4AC determines is not shared. Similarly, for privilege subsumption, a true positive is a privilege of the second NLACM that is subsumed by that of the first, and this was detected by LLM4AC’s differencing procedure. A true negative is a privilege not subsumed by that of the first that was not detected by LLM4AC’s differencing procedure. A false positive is a privilege not subsumed that LLM4AC’s differencing procedure incorrectly determines is subsumed. A false negative is a subsumed pair of privileges that LLM4AC determines is not.
Differencer Benchmark Details. To evaluate LLM4AC with respect to policy auditing and comparison, we construct three NLACMs: a base NLACM containing NL, a SQL NLACM containing the ground truth SQL in the base, and a perturbed NLACM, which is the base NLACM with a NL perturbation applied, such that the resulting NL is syntactically different, but semantically equivalent (e.g., all column names are replaced with synonyms). Then, policy auditing is differencing the base NLACM to the SQL NLACM. Policy comparison is the accuracy of differencing the base and perturbed NLACMs.
An IBAC-DB system needs to have high cross-domain performance–that is, the differencer should accurately compare policies on databases with a variety of schemas and policy specifications unseen by the model. To test this using our benchmark, we construct these NLACMs from: (i) databases spanning multiple training domains, and (ii) syntactically different but semantically equivalent NL (e.g., column names expressed using carrier phrases instead of their actual names). We gather this data from an existing benchmark, Dr. Spider (Chang et al., 2023), a NL2SQL benchmark with databases spanning several training domains, and several NL perturbations.
There are only 5 databases in Dr. Spider that have NL-query-perturbation triples for all perturbation types. We create the equivalent base-SQL-perturbed NLACMs for each of these databases. We list the databases and some basic characteristics here:
@@
NLACM Construction. To construct the necessary NLACM pairs for policy comparison and auditing, we extract the NL describing database views from Dr. Spider. We automatically generate the roles and privileges in the NLACM. To experiment with NL perturbations of roles and views as well, we define and use the following perturbations:
@@ 5.1.2. The Synthesizer Benchmark
Synthesizer Evaluation Metric. We input this NLACM to LLM4AC’s Synthesizer and test the execution accuracy of the result. Specifically, we execute the resulting GRANT statements on one copy of the databases, and execute the ground truth on a separate copy of the databases. PostgreSQL stores the results as a list of privileges in a table. We compare these tables to see if they are equal.
Synthesizer Benchmark Details. To gather these queries of varying types and complexity, and their NL to be translated into SQL, we use a mix of the TPC-H queries and queries from several Spider databases.
@@ 5.2. LLM4AC Differencing: Macrobenchmark
In this section, we use our IBAC-DB benchmark to evaluate the accuracy of LLM4AC’s differencer.
Evaluation Baselines. We compare LLM4AC’s prompting differencing procedure to the following baselines for role-view mapping:
@@
    Plain LLM: LLM4AC prunes out incorrect answers by only including candidates that share DB literals. But is this DB literals comparison actually necessary? To explore this, we include a plain use of the LLM where answers are not pruned out. Everything else is kept the same.
@@
For all baselines except Jaccard, we adjust them for comparing NL and SQL by generating a NL description of the SQL query, and comparing the NL description to a list of NL roles/views of the other NLACM.
@@
Figure 6. NL vs NL Role-View Mapping F1 Errors across all DBs and Perturbations. Lower is better.
Figure 7. NL vs NL Privilege Subsumption F1 Errors across all DBs and Perturbations. Lower is better.
Policy Comparison.: Figure 7 and Figure 6 shows that, across databases and perturbations, role-view mapping appears to be more difficult than privilege subsumption. Further, LLM4AC or sentence embeddings outperform other role-view mapping methods.
Figure 8. NL vs SQL Role-View Mapping F1 Errors across all DBs and Perturbations. Lower is better.
Figure 9. NL vs SQL Privilege Subsumption F1 Errors across all DBs and Perturbations. Lower is better.
Implementation Comparison.: Figure 9 and Figure 8 show that overall, LLM4AC outperforms the baselines. For other methods, the F1 errors are all higher compared to those of policy comparison. The F1 error for Plain LLM is higher than for other methods, which initially suggests that using LLMs is not preferable. However, this is because all the errors are still due to mistakes that could have been fixed by comparing DB literals. In this case, this would simply be a matter of using SQL parsing and then string matching. LLM4AC incorporates this observation, leading to its high performance. The other methods have high errors, likely due to the noise introduced when translating the SQL to NL. For example, the view ChatGPT explains the view definition CREATE VIEW query2view0 SELECT Model FROM CAR_NAMES GROUP BY Model ORDER BY count(*) DESC LIMIT 1 using NL: This SQL query creates a view called ”query2view0” that selects the ”Model” column from a table called ”CAR_NAMES”. It then groups the results by ”Model”, counts the number of occurrences of each model, and orders the results in descending order based on the count. Finally, it limits the output to only show the model with the highest count. Essentially, this query finds the most common car model in the ”CAR_NAMES” table.. The baselines must somehow match this long description to the NL question: How many car models are produced by each maker ? Only list the count and the maker full name .
Overall, we observe that our baselines perform quite well at policy comparison. Even Jaccard similarity has a F1 of 0.7 on the Dr. Spider perturbations. This is because many of these perturbations do not change many tokens in the original text, only replacing certain phrases (e.g., column names or DB values) with synonyms, phrases, etc. This also helps explain why word embeddings can be even worse than Jaccard similarity. This is because by naively averaging embeddings over words, it is no longer possible to match on DB literals, which would enable more accurate matching.
@@ 5.3. LLM4AC Synthesizer: Macrobenchmark
Table 1. LLM4AC Synthesizer Execution Accuracy. The only errors were nested queries and multi-way joins.
We see that LLM4AC’s synthesizer is mostly accurate, failing at multi-way joins and nested queries. We observe these errors for the following reasons: (1) C3 can occasionally choose the wrong tables for a multi-way join (2) C3 can “forget” that a nested query is needed.
@@ 5.4. LLM4AC Differencer: Microbenchmark
Figure 10. NL vs NL Role-View Mapping F1s by Perturbation. Darker is better, and shading is row-wise.
Figure 11. NL vs NL Privilege Subsumption F1s by Perturbation. Darker is better, and shading is row-wise.
Figure 12. NL vs NL Role-View Mapping F1s by DB. Darker is better, and shading is row-wise.
Figure 13. NL vs NL Privilege Subsumption F1s by DB. Darker is better, and shading is row-wise.
@@ 5.4.1. LLM4AC vs Baselines, by DB/Perturbation
Policy Comparison. When we compute the F1s for each perturbation, we find that LLM4AC vastly outperforms all other methods, followed by sentence embeddings (Figure 10 and Figure 11). We find that this is because naively providing all inputs to prompting generates incorrect answers that are fixed simply by comparing sentences with shared DB literals instead. On the other hand, sentence embeddings have more obscure errors: namely, when the correct answer is a more abstract sentence (e.g., ”How many orchestras has a conductor done?”) sentence embedding similarity is incorrect. We find the same pattern over databases (Figure 12 and Figure 13).
Figure 14. NL vs SQL Role-View Mapping F1s, by Perturbation. Darker is better, and shading is row-wise.
Figure 15. NL vs SQL Privilege Subsumption F1s by Perturbation. Darker is better, and shading is row-wise.
Figure 16. NL vs SQL Role-View Mapping F1s by DB. Darker is better, and shading is row-wise.
Figure 17. NL vs SQL Privilege Subsumption F1s by DB. Darker is better, and shading is row-wise.
Implementation Comparison. Figure 14 and Figure 15 show the results for differencing stratified by perturbation. Based on the F1 scores listed, we do not find a difference between the performance of LLM4AC relative to other role-view mapping methods with respect to perturbation. Similarly, Figure 16 and Figure 17 show the results stratified by database. We do not find a difference between the performance of LLM4AC relative to other role-view mapping methods with respect to database, either.

### v2

@@ 1. Introduction
Over the last several years, there has been a proliferation of easy-to-use software tools for data science. Organizations of all sizes and from all industries are applying data-driven models to make decisions. This trend, which is often called the democratization of data science (Chou et al., 2014), is as worrying as it is exciting. On one hand, some of humanity’s most difficult scientific challenges are being addressed with data science from particle physics to drug discovery. On the other hand, there are numerous cautionary tales of data misuse, data breaches, data quality issues, and improper data sharing (Monroe-White et al., 2021; Janssen et al., 2020; Kerber, 2020; Seh et al., 2020; Fahey, 2021). Configuring database deployments, to manage availability, usability, integrity, and security of the data in enterprise systems, has consequently emerged as a key discipline in enterprise data management. Access control is one of the most important technical challenges when configuring databases, where a system must ensure that only authorized users can access particular data assets.
In any sizeable organization, it is impractical to exhaustively map individual users to data assets. Thus, role-based access control (RBAC) was a key innovation and is widely supported by data management platforms where groups of users can be selectively given access (Sandhu et al., [n.d.]). Such forms of access control are well-studied in enterprise SQL databases, where roles, views, and GRANT statements are already used to restrict user access to data (Chaudhuri et al., 2011; Bertino et al., 2011; Bertino and Sandhu, 2005). Similarly, many databases additionally support discretionary access control (DAC) where roles can be given delegation privileges. RBAC and DAC help reduce the administrative burden in defining access control rules for a database.
@@
  • LLM4AC for Access Control Privileges: A system built from a reference architecture for an IBAC-DB implementation that provides a novel format for specifying access control policies, automatically implementing them, and comparing implementations to policies for correctness.
@@
  • Generalizing to Other Deployment Requirements: we propose two other systems that generalize LLM4AC to other requirements needed to deploy databases: one requirement-specific system that synthesizes and differences implementation using LLM prompts specialized to multiple types of deployment requirements, such as other access control requirements (e.g., role hierarchies, temporal access control), and one generalized system that uses user-inputted requirement types to analyze the access control policy for these requirement types, and synthesize and difference the corresponding access control implementation.
@@
This paper is structured as follows: in Section 2, we describe the key definitions, problem statement, and related work describing synthesis and auditing of access control. In Section 3, we describe a reference architecture specifically for implementing access control privileges (permitted SQL operators to roles on database views). In Section 4, we describe our system, LLM4AC, which is an instantiation of the reference architecture using ChatGPT and Postgres. In Section 5, we evaluate LLM4AC. In Sections 6 and 7, we describe and evaluate our requirement-specific system design respectively. Similarly, in Sections 8 and 9, we describe and evaluate our generalized system design. We conclude in Section 10.
@@ 2.4. Existing NL2SQL Capabilities
Based on these examples and initial experiments, rather than trying to train NL2SQL systems, we hypothesize if access control policies can be formatted to clearly reflect a role, privilege, and view, and we can correctly represent access control policies in terms of the database roles, views, and privileges that correspond to the policy, then NL2SQL methods adapted for access control can correctly implement access control policies.
Why Trust LLMs for Access Control? There are salient differences between the traditional NL2SQL problem for generating SQL DML, and the NL2SQL problem for DDL (including access control). NL ambiguity for current NL2SQL has been observed to lead to multiple correct answers (Floratou et al., 2024) but access control has only one correct implementation (only one set of GRANT statements can correctly implement a database access control policy). Considering this, and the very limited (ideally non-existent) scope for error in access control implementation, why should we trust LLMs for access control? In this paper, we do not assume the output of a NL2DDL system will directly be used by an organization. Instead, we propose a system that will make access control implementation less tedious and error-prone if used in the following workflow: (1) Users will write access control policies (2) The access control policies will be inputted to the system (3) The system will synthesize SQL, and audit the synthesized implementation with respect to the policy (4) The end user will receive the implementation and audit report, which they will use to correct implementation errors. The ease and reliability with which the end user can correct the output implementation will depend on the system’s synthesis and auditing accuracy.
This workflow requires accurate methods of evaluating NL2SQL systems for DDL. NL2SQL systems for DML are evaluated via execution accuracy by comparing outputted recordsets to one or more ground truths. On the other hand, executing access control SQL only changes the database state, and returns nothing. Therefore, a NL2SQL system for access control would currently have to be evaluated by: (i) manually devising the only ground truth implementation, and then (ii) combinatorially generating tests to compare whether allowed operations under the ground truth are allowed under the system-generated implementation, and vice versa (Bertolino et al., 2018; Li et al., 2015; Zhang et al., 2023). These steps are tedious and error-prone. This is why we propose auditing implementation by using LLMs to directly compare the NL policy to its SQL implementation.
@@ 4.2. The Differencer, In-Depth
LLM4AC At Scale. LLM4AC scalably maps roles and views by using prompts that match single roles/views from $M_{1}$ to lists of roles/views $M_{2}$ (i.e., perform a nearest-neighbor-if-exists on roles/views of $M_{2}$). For example, ChatGPT has a context window length of 8192 tokens, and a list of 10 view definitions, each roughly 50 tokens long, on average, resulting in only 10 calls of 500 tokens each, with 500 more tokens of output. This costs 5500 tokens. This is cheaper compared to 100 pairwise comparisons, each consisting of at least 100 tokens and 1 token of output (a yes/no token) (10001 tokens). This enables cheaper differencing of large NLACMs, such as comparing a policy to a production database.
5. LLM4AC Evaluation
In this section, we propose a benchmark for evaluating IBAC-DB systems (Section 5.1). We use this benchmark to show that our system, LLM4AC, can accurately synthesize and difference access control policies (Sections 5.2, 5.3 and 5.4), especially compared to other baselines. Finally, we show that LLM4AC performs accurately when external knowledge is needed using a real-world access control policy (Section 5.5). For all experiments, the DBMS is PostgreSQL. We use ChatGPT (gpt-3.5-turbo) as the LLM, and ChatGPT’s APIs for all prompting.
@@ 5.1. IBAC-DB Benchmark Design
We design a benchmark to test IBAC-DB systems that implement access control policies in PostgresQL. The goal of the benchmark is to test how accurately an IBAC-DB system synthesizes a SQL implementation of a NLACM policy, and differences two NLACMs.
@@ 5.1.1. The Differencer Benchmark
Differencer Evaluation Scenarios. We measure differencing with respect to two scenarios: (i) policy auditing: compare a NLACM containing NL to its implementation on the DB, which is a NLACM containing SQL. (ii) policy comparison: compare two NLACMs containing NL.
Differencer Evaluation Metric. We evaluate differencing with respect to F1 scores, where a positive indicates that both NLACMs share a privilege. We also use F1 scores to evaluate role-view mapping and privilege subsumption. A role-view mapping positive indicates a role-view pair is shared by both NLACMs. A privilege subsumption positive indicates the privileges of the second NLACM are subsumed by the first. In cases where performance is near-perfect, we instead use the F1 error (1 - F1) for presentation.
Differencer Benchmark Details. To evaluate LLM4AC with respect to policy auditing and comparison, we construct three NLACMs: a base NLACM containing NL, a SQL NLACM containing the ground truth SQL, and a perturbed NLACM, which is the base NLACM with a NL perturbation applied, such that the resulting NL is syntactically different, but semantically equivalent (e.g., all column names are replaced with synonyms). Then, policy auditing is differencing the base NLACM to the SQL NLACM. Policy comparison is differencing the base and perturbed NLACMs.
An IBAC-DB system needs to have high cross-domain performance–that is, the differencer should accurately compare policies on databases with a variety of schemas and policy specifications. To test this using our benchmark, we construct these NLACMs from: (i) databases spanning multiple training domains, and (ii) syntactically different but semantically equivalent NL (e.g., column names expressed using carrier phrases instead of their actual names). We gather this data from an existing benchmark, Dr. Spider (Chang et al., 2023), a NL2SQL benchmark with databases spanning several training domains, and several NL perturbations.
There are only 5 databases in Dr. Spider that have NL-query-perturbation triples for all perturbation types. We create the equivalent base, SQL, and perturbed NLACMs for each of these databases. We list the databases and some basic characteristics here:
@@
NLACM Construction. To construct NLACM pairs for policy comparison and auditing, we extract the NL describing database views from Dr. Spider. We automatically generate the roles and privileges in the NLACM. To experiment with NL perturbations of roles and privileges as well, we define and use the following perturbations:
@@ 5.1.2. The Synthesizer Benchmark
Synthesizer Evaluation Metric. We input this NLACM to LLM4AC’s Synthesizer and test the accuracy of the result by executing the resulting GRANT statements on one copy of the databases, and executing the ground truth on a separate copy of the databases. PostgreSQL stores the results as a list of privileges in a table. We compare these tables to see if they are equal.
Synthesizer Benchmark Details. To gather the NL and SQL of these queries of varying types and complexity, we use a mix of the TPC-H queries and queries from several Spider databases.
@@ 5.2. LLM4AC Differencing: Macrobenchmark
Evaluation Baselines. We compare LLM4AC’s differencing procedure to the following baselines for role-view mapping:
@@
    Plain LLM: LLM4AC prunes out incorrect answers by only including candidates that share DB literals. We quantify the benefits of literal pruning by including a LLM baseline where answers are not pruned out. Everything else is kept the same.
@@
For all baselines except Jaccard, we adjust them to compare NL and SQL by generating a NL description of the SQL query, and comparing it to a list of NL roles/views of the other NLACM.
@@
Plain LLM Sentence Word Jaccard LLM4AC
NL vs NL
Total
RV 1-F1
0.149 0.084 0.220 0.204 0.041
NL vs NL
Total
Priv 1-F1
0.119 0.116 0.122 0.116 0.003
NL vs SQL
Total
RV 1-F1
0.553 0.765 0.632 0.837 0.035
NL vs SQL
Total
Priv 1-F1
0.021 0.007 0.014 0.014 0.006
Table 1. LLM4AC Differencer F1 Errors across all DBs and Perturbations. Lower is better.
Table 1 shows total F1 errors calculated over all DBs and perturbations for privilege subsumption and role-view mapping. The first two rows are policy comparison results (NL vs NL), and the second two rows are policy auditing (NL vs SQL).
Policy Comparison.: The first two rows of Table 1 show that, across databases and perturbations, role-view mapping seems more difficult than privilege subsumption. Further, LLM4AC or sentence embeddings outperform other role-view mapping methods.
Policy Auditing.: The last two rows of Table 1 show that overall, LLM4AC outperforms the baselines. For other methods, the F1 errors are all higher compared to those of policy comparison. The F1 error for Plain LLM is generally higher than for other methods, which initially suggests that using LLMs is not preferable. However, this is because all the errors are still due to mistakes that could have been fixed by comparing DB literals. LLM4AC incorporates this observation, leading to its high performance.
@@ 5.3. LLM4AC Synthesizer: Macrobenchmark
Table 2. LLM4AC Synthesizer Accuracy. The only errors were nested queries and multi-way joins.
Table 2 shows that LLM4AC’s synthesizer is mostly accurate, failing at multi-way joins and nested queries. This is because: (1) C3 can occasionally choose the wrong tables for a multi-way join (2) C3 can “forget” that a nested query is needed.
@@ 5.4.1. LLM4AC vs Baselines, by DB/Perturbation
Policy Comparison. When we compute the F1s for each perturbation, we find that LLM4AC vastly outperforms all other methods, followed by sentence embeddings (Figure 6, (A)). This is because naively providing all inputs to prompting generates incorrect answers that are fixed by comparing sentences with shared DB literals instead. On the other hand, sentence embeddings have more obscure errors: namely, when the correct answer is a more abstract sentence (e.g., ”How many orchestras has a conductor done?”) sentence embedding similarity is incorrect. We find the same pattern when stratifying over databases.
Policy Auditing. (Figure 6, (B)) show the results for differencing stratified by perturbation. We do not find a difference between LLM4AC and other role-view mapping methods with respect to perturbation or database.